If you have two flag poles in front of the court house 12ft tall. The distance from the top of one pole to the base of the other is 20ft . What is the distance between the two flag pole?
Answer by edjones(8007)

a=12, c=20
a^2+b^2=c^2
12^2+b^2=20^2
144+b^2=400
b^2=400-144
b^2=256
b=16' sqrt of each side
.
Ed